OpenStudy (anonymous):

4x-1=4x^2 find the solution, step by step please.

OpenStudy (anonymous):

\[4x ^{2}-4x+1=0\] split the middle term.. so you have \[4x ^{2}-2x-2x+1=0\] then factor them separately.. \[2x(2x-1)-(2x-1)\] so then the answer is: \[(2x-1)(2x-1) = 0\] all you have to do now is equate (2x-1) to 0. \[2x-1=0\]\[2x=1\]\[x=1/2\]

OpenStudy (anonymous):

oh im sorry i forgot to put the negetive in front of the 4x^2